# Decoherence and Probability

### v2

@@ 4.3. Emergent Probability
Full articulation of our account of the emergence probability in the context of classical and quantum possibility space models will require appeal to a cluster of conceptual innovations articulated in the philosophical literature over the last decades.
Most foundational, is the account of emergence due to Butterfield (2011). On Butterfield’s account, emergence should be understood to indicate the deduction of a novel and robust behaviour. Furthermore, at least in a weak sense, such emergence can obtain on the way to the limit that forms part of the relevant deduction. Butterfield’s account of emergence was subsequently refined and further articulated in a manner well-suited to our purposes by Palacios (2018, 2019, 2022). Of particular relevance the notion of coarse-grained emergence where a coarse-grained description of a system emerges from a fine-grained description, if and only if the former has terms denoting properties or behaviour that are novel and robust with respect to the latter (Palacios 2022, p.). Also important will be the idea of combined dual synchronic and diachronic derivation of an emergent behaviour (Palacios 2022, p.).
@@
Finally, and crucially, our account of the emergence of probability will also require us to draw upon the idea of a ‘factual’ interpretation of the semi-classical limit due to Feintzeig (2020). According to the account of Feintzeig (2020), which extends ideas of Rohrlich (1989) and Fletcher (2019), we should understand the limiting procedure an approximation at certain scales with the change in the scale is determined by the error bounds and units set by a measurement procedure that is used to probe the system. This approach allows us to recover the intuitive notion that the limit can be understood in terms of ‘zooming out’ from the characteristic quantum scale set by , by caring less and less about the microscopic details.
@@
The requirement for this particular approach to the semi-classical limit can be most clearly articulated in the context of the contrasting treatment we will adopt towards the quantities in the theory and the Wigner function. Not only is the semi-classical limit of the Wigner function not always well-behaved but for various important classless of quantum states, in particular coherent states, we find that as , the Wigner function approximates a -function (Berry 1977; Curtright et al. 2013; Mariño 2021). In such circumstances, consistently applying a ‘counterfactual’ account of the semi-classical limit, where we consider a physical world in which physical dynamics is constituted by the limit behaviour, would seem to allow us only to recover the deterministic subset of classical possibility space models.
@@
By contrast, on the factual approach, there is no contradiction in ‘zooming out’ such that terms are neglected and yet retaining non-trivial spread on the Wigner function even for coherent states. Ultimately, this is to focus on behaviour that, in Butterfield’s terms, emerges ‘on the way to the limit’, in the sense that so long as the scale of the measurement probe is appropriately represented by error bounds and units so that is far enough on the way to the limit, then one is justified in ignoring the terms precisely because they will be smaller than what is detectable at that scale.131313Thanks to Ben Feintzeig for helpful discussion on this point This is precisely the intuition one has in considering Figure 1 and the role of the box of area .
@@
Our proposal is then to deploy a combined Butterfield-Palacios-Feintzeig approach to coarse-grained emergence of semi-classical behaviour under a factual interpretation of the semi-classical limit taken as an approximation. Recall, that a coarse-grained description of a system emerges from a fine-grained description, if and only if the former has terms denoting properties or behaviour that are novel and robust with respect to the latter. In our case the ‘fine-grained’ description is the full quantum phase space model and the ‘coarse-grained’ description is the semi-classical phase space model which is such that the expectation values and expressions truncated are isomorphic to a classical phase space model.
@@
We find emergence in the sense of Palacios (2022) account of coarse-grained emergence specifically since we have: (i) a fine-grained/coarse-grained distinction picked out by phase space areas at order /at order much bigger than ; (ii) the coarse-grained description has features that are not features of the fine-grained description, specifically conservation and localisation of the (quasi)-probability density; (iii) the behaviour represented by the fine-grained description exists at the same time as the behaviour represented by the coarse-grained description (i.e. we have synchronic emergence); (iv) the coarse-grained description refers to some behaviour that is insensitive to variation of the microphysical details that characterise a particular token (i.e. we have robustness); (v) the coarse-grained level depends on the fine-grained level in the sense that every change in the coarse-grained level must imply a change in the fine-grained level (i.e. we have supervenience).
@@
Putting everything together. Quasi-probabilistic emergence via decoherence consists in the derivation of the novel and robust behaviour of Wigner positivity at finite values of the relevant parameters (time and temperature in realistic models). Coarse-grained emergence via a factual interpretation of the semi-classical limit consists in the derivation of the behaviours of conservation and localisation in the coarse-grained description that are novel and robust with respect to the fine-grained description. Whereas the quasi-probabilistic emergence via decoherence is dependant on the decoherence time scale and is thus ‘diachronic’. By contrast, the semi-classical emergence is dependent upon phase space areas in units of and is thus ‘synchronic’. A schematic diagram for the relevant pattern of interrelations is provided in Figure 2. The combined double emergence picture is sufficient to deduce a classical possibility space model from a quantum possibility space model. Classical probability in phase space is emergent from an underlying quantum phase space dynamics.
@@
We take ourselves to have comprehensive formal and physical account of the emergence of probability from quantum theory. This is not, however, an interpretation of the theory or a solution to the measurement problem. Rather, our aim has been to clarify the sense in which principles of quantum mechanics require generalised probabilistic concepts and how such concepts and how such generalised can be connected to classical probability via physical arguments. In the following section, we will summarise our main arguments and offer an outlook regarding the further work that would be required to offer an response to the interpretational issue.
@@ 5. Recapitulation and Outlook
Let us return to the original dialectic with which we started our analysis of probably and decoherence. Recall, in particular, that the arguments of Dawid and Thébault (2015) that a certain package of interpretative moves concerning probability and the quantum formalism leads to an incoherent conclusion. The foregoing analysis allows us to consider the contraposition this argument. That is, we have now established a plausible framework for the analysis of classical and quantum probability within which any coherent interpretation must be expected to operate.
@@
Second, there is an interesting connection between our analysis, proposal for quantum measure theory (Sorkin 2010; Clements et al. 2017) and generalised probabilistic structures found within the decoherent histories programme (Gell-Mann and Hartle 1996; Halliwell 2010). It can be proved that the diagonal elements of the decoherence functional are equivalent to a ‘quantal-measure’ which is a specific form of our quasi-measure that obeys a particular (non-classical) sum rule on the algebra of events (Sorkin 1994; Dowker and Wilkes 2022). The decoherent histories framework thus is a partial interpretation of a quasi-probability structure in precisely our terms. Since there is an explicit dependence on coarse-graining in this approach to the emergence of classical probability there is plausible path for both reconstructing our analysis in histories terms and applying such a quantum measure theory version of the programme as a full interpretation of the theory.
@@
Fourth, much more could be said about the interaction between diachronic and synchronic limits the context of classical and quantum probability. For example, it would be of significant physical and philosophical interest to more fully understand the relation between the semi-classical and long time limits. It was recently demonstrated by Bonds et al. (2024), that application of the method of arbitrary functions allows one to explicitly derive a classical possibility model in the case of the quantum oscillator via the combined and limits. Furthermore, this approach allows for the recovery of the Born rule when combined with a toy model of quantum measurement in terms of a perturbation treated as a random variable with (almost) arbitrary initial probability distribution in a manner inspired by the ‘flea’ perturbation of Landsman and Reuvers (2013). This approach would offer a further alternative strategy for extending our analysis towards a full interpretation of the theory.141414Relatedly, it would also be interesting to consider the connection between our framework and recent work on the semi-classical limit and the emergence of classical probability in the context of open systems due to Layton and Oppenheim (2023) and Hernández et al. (2023).
@@ Acknowledgements
Work on this paper has drawn on a large number of discussions with a wide variety of helpful colleagues over the last few years. We are particularly grateful to Simon Saunders and Alex Franklin for continued highly enjoyable interactions and to Ben Feintzeig for constructive critical engagement. Further thanks are also due to Jim Al-Khalili, Kevin Blackwell, Dorje Brody, Adam Caulton, Fay Dowker, Sam Fletcher, Gui Franzmann, Simon Friedrich, Sean Gryb, Jonte Hance, Nick Huggett, Eddy Keming Chen, Jason Konek, James Ladyman, Tushar Menon, Richard Pettigrew, James Read, Christian Röken, Tony Short, Jer Steeger, Cristi Stoica, and Lev Vaidman, and two anonymous referees (apologies if we have missed anyone out). Thanks also to audiences in Surrey, Bonn, Oxford, and Bristol. We are also appreciative to Cosmas Zachos for helpful written comments on a draft manuscript. Work on this paper was supported by the John Templeton Foundation grant number 62210. The views expressed may not represent the opinions of the Foundation. This research was in part funded by the Swedish Research Council grant number 2022-01893_VR.

### v3

@@ 4.3. Emergent Probability
Our account of the emergence probability in the context of classical and quantum possibility space models will rely upon appeal to a cluster of conceptual innovations from the philosophical literature over the last decades.
Most foundational is the account of emergence due to Butterfield (2011). On Butterfield’s account, emergence should be understood to indicate the existence of novel and robust behaviour which can be deduced, within a model of that behaviour, by taking an appropriate mathematical limit (this is Butterfield’s 1:Deduce). Furthermore, at least in a weaker (i.e. approximate) sense, such emergence can obtain on the way to the limit that forms part of the relevant deduction (this is Butterfield’s 2:Before). Butterfield’s account of emergence was subsequently extended in terms of a notion of coarse-grained emergence by Palacios (2018, 2019, 2022). We will provide details of that account shortly.
@@
Next, and crucially, our account of the emergence of probability will also require us to draw upon the idea of a ‘factual’ interpretation of the semi-classical limit due to Feintzeig (2020). According to the account of Feintzeig (2020), which extends ideas of Rohrlich (1989) and Fletcher (2019), we should understand the limiting procedure as an approximation at certain scales with the change in the scale determined by the error bounds and units set by a measurement procedure that is used to probe the system. The factual approach allows us to recover the intuitive notion that the limit can be understood in terms of ‘zooming out’ from the characteristic quantum scale set by by caring less and less about the microscopic details. The contrast is with a ‘counterfactual’ approach to the limiting procedure in which we consider the physics of possibilities apart from the actual world in which Planck’s constant takes a different value that in the limit goes to zero.
@@
The requirement for the factual approach to the semi-classical limit can be explained most clearly in the context of the contrasting treatment we will adopt towards the quantities in the theory and the Wigner function. The semi-classical limit of the Wigner function is not always well-behaved (Berry 1977). Furthermore, for various important classes of quantum states, in particular coherent states, we find that as , the Wigner function approximates a -function (Berry 1977; Curtright et al. 2013; Mariño 2021). In such circumstances, applying a ‘counterfactual’ account of the semi-classical limit, where we consider a physical world in which physical dynamics is constituted by the limit behaviour, allows us only to recover the deterministic subset of classical possibility space models since, for consistency, the interpretation requires us to both neglect the terms and work in the limit the probability density is focused on a single point.
@@
By contrast, on the factual approach, since we have different interpretational resources, we can consistently ‘zoom out’ such that terms are neglected and yet retain non-trivial spread on the Wigner function, even for coherent states. Ultimately, this is to focus on behaviour that, in Butterfield’s terms, emerges ‘on the way to the limit’, in the sense that so long as the scale (determined by the error bounds and units set by a measurement procedure) is far enough on the way to the limiting case (in which microscopic details can be neglected entirely), then one is justified in ignoring the terms precisely because they will be smaller than what is detectable at that scale. This is precisely the intuition one has in considering Figure 1 and the role of the box of area .
@@
Our proposal is then to deploy a combined Butterfield-Palacios-Feintzeig approach to coarse-grained emergence of semi-classical behaviour under a factual interpretation of the semi-classical limit taken as an approximation. According to (Palacios 2022, p.39), a coarse-grained description of a system emerges from a fine-grained description, if and only if the former has terms denoting properties or behaviour that are novel and robust with respect to the latter. In our case, the ‘fine-grained’ description is the full quantum phase space model and the ‘coarse-grained’ description is the semi-classical phase space model which is such that the expectation values and expressions truncated are isomorphic to a classical phase space model.
@@
We find emergence in the sense of Palacios (2022) account of coarse-grained emergence specifically since we have: (i) a fine-grained/coarse-grained distinction picked out by phase space areas at order /at order much bigger than ; (ii) the coarse-grained description has features that are not features of the fine-grained description, specifically local conservation and localisation of the (quasi)-probability density; (iii) the behaviour represented by the fine-grained description exists at the same time as the behaviour represented by the coarse-grained description (i.e. we have synchronic emergence); (iv) the coarse-grained description refers to behaviours (local conservation and localisation) that are insensitive to (robust under) variation of the microphysical details that characterise a particular token (for example, with respect to variation of fine-grained details ); (v) the coarse-grained level depends on the fine-grained level in the sense that every change in the coarse-grained level must imply a change in the fine-grained level (i.e. we have supervenience).
@@
Quasi-probabilistic emergence via decoherence consists in the derivation of the novel and robust behaviour of Wigner positivity at finite values of the relevant parameters (time and temperature in realistic models). Coarse-grained emergence via a factual interpretation of the semi-classical limit consists in the derivation of the behaviours of conservation and localisation in the coarse-grained description that are novel and robust with respect to the fine-grained description. The quasi-probabilistic emergence via decoherence is dependant on the decoherence time scale and is thus ‘diachronic’. By contrast, the semi-classical emergence is dependent upon phase space areas in units of and is thus ‘synchronic’. A schematic diagram for the relevant pattern of interrelations is provided in Figure 2. The combined double emergence picture is sufficient to deduce a classical possibility space model from a quantum possibility space model. Classical probability in phase space is emergent from an underlying quantum phase space dynamics.
@@
We take ourselves to have provided a comprehensive formal and physical account of the emergence of probability from quantum theory. This is not, however, an interpretation of the theory or a solution to the measurement problem. Rather, our aim has been to clarify the sense in which principles of quantum mechanics require generalised probabilistic concepts and how such concepts and how such generalised can be connected to classical probability via physical arguments. In the following section, we will summarise our main arguments and offer an outlook regarding the further work that would be required to offer a response to the interpretational issue.
@@ 5. Recapitulation and Outlook
Let us return to the original dialectic with which we started our analysis of probability and decoherence. Recall, in particular, that the arguments of Dawid and Thébault (2015) were that a certain package of interpretative moves concerning probability and the quantum formalism leads to an incoherent conclusion. The foregoing analysis allows us to consider the contraposition this argument. That is, we have sought to establish a framework for the analysis of classical and quantum probability within which any coherent interpretation must be expected to operate.
@@
Second, there is an interesting connection between our analysis, proposals for quantum measure theory (Sorkin 2010; Clements et al. 2017), and generalised probabilistic structures found within the decoherent histories programme (Gell-Mann and Hartle 1996; Halliwell 2010). In particular, it can be proved that the diagonal elements of the decoherence functional are equivalent to a ‘quantal-measure’ which is a specific form of our quasi-measure that obeys a particular (non-classical) sum rule on the algebra of events (Sorkin 1994; Dowker and Wilkes 2022). The decoherent histories framework thus is a partial interpretation of a quasi-probability structure in precisely our terms since it involves the weighting of quantum possibilities via the decoherence functional. Since there is an explicit dependence on coarse-graining in this approach to the emergence of classical probability there is a plausible path for both reconstructing our analysis in histories terms and applying such a quantum measure theory version of the programme as a full interpretation of the theory.
@@
Fourth, much more could be said about the interaction between diachronic and synchronic limits in the context of classical and quantum probability. For example, it would be of significant physical and philosophical interest to more fully understand the relation between the semi-classical and long time limits. It was recently demonstrated by Bonds et al. (2024), that application of the method of arbitrary functions allows one to explicitly derive a classical possibility model in the case of the quantum oscillator via the combined and limits. Furthermore, this approach allows for the recovery of the Born rule when combined with a toy model of quantum measurement in terms of a perturbation treated as a random variable with (almost) arbitrary initial probability distribution in a manner inspired by the ‘flea’ perturbation of Landsman and Reuvers (2013). This approach would offer a further alternative strategy for extending our analysis towards a full interpretation of the theory.111111Relatedly, it would also be interesting to consider the relation to the work of Layton and Oppenheim (2023) and Hernández et al. (2023).
@@ Acknowledgements
Work on this paper has drawn on a large number of discussions with a wide variety of helpful colleagues over the last few years. We are particularly grateful to Simon Saunders and Alex Franklin for continued highly enjoyable interactions and to Ben Feintzeig for constructive critical engagement. Further thanks are also due to Jim Al-Khalili, Kevin Blackwell, Dorje Brody, Adam Caulton, Fay Dowker, Sam Fletcher, Gui Franzmann, Simon Friedrich, Sean Gryb, Jonte Hance, Nick Huggett, Eddy Keming Chen, Jason Konek, James Ladyman, Tushar Menon, Richard Pettigrew, James Read, Christian Röken, Tony Short, Jer Steeger, Mritunjay Tyagi, Cristi Stoica, and Lev Vaidman, and three anonymous referees (apologies if we have missed anyone out). Thanks also to audiences in Surrey, Bonn, Oxford, and Bristol. We are also appreciative to Cosmas Zachos for helpful written comments on a draft manuscript. Work on this paper was supported by the John Templeton Foundation grant number 62210. The views expressed may not represent the opinions of the Foundation. This research was in part funded by the Swedish Research Council grant number 2022-01893_VR.